[Brahms-l] Shift report 20050219 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 19 2005 - 16:01:43 EST
Shifter leader was CN
Shifters were 


1. once the beam was up it ran smoothly.

2.  Runs:        13266,  13267, 13294 

    FS@4B3450:    141k,  149k, 288k
    MRS@40A1050:  270k,  241k, 191k
    begun r13295

non-tpc ped run: 13291
tpc ped run: 13293

3. Problems with Tpc1, however apparently working now.    


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 19 16:01:59 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 19 2005 - 16:02:07 EST